# EPiC: Cost-effective Search-based Prompt Engineering of LLMs for Code Generation

### v1

@@ 4.3 Evaluation Metrics
To evaluate the quality of the generated code by LLMs, we utilized the pass@k metric. This metric, introduced in [8], is advantageous over metrics such as CodeBLEU and ROUGE because it evaluates code behavior rather than text semantics, which are the focus of the latter metrics. The formula for pass@k is presented in Equation 1. In this equation, $E$ denotes the expected value over the problems, $n$ is the total number of generated samples per task (code samples) and $c$ is the number of correct samples. For a single problem, this metric estimates the probability that at least one of top $k$ samples is correct. When applied to multiple problems, it evaluates the expectation across all problems.
@@
When comparing EPiC with baselines, we also report Mann-Whitney U-test results, which is a non-parametric significant test, to consider the randomness introduced by our algorithm. The tests compare the distribution of 10 pass@1 results for EPiC (using 10 random seeds) with the pass@1 of each baseline, separately. Note that here we only rerun EPiC with 10 random seeds to consider its internal randomness introduced in the algorithm and do not rerun the baselines as they do not have internal randomness. In other words, we do not study the potential randomness of LLMs results and their effect on each baseline, as this is not reported in the original papers and would go beyond the scope of this paper.
4.4 Objective Function
We use $\mathcal{X}$ and $\mathcal{Z}$ to denote the prompt space and output responses respectively. Let $S:\mathcal{X}\rightarrow\mathcal{Z}$ be the LLM that takes an input $x\in\mathcal{X}$ and outputs a response $z\in\mathcal{Z}$. The output response is the code implementation of the given prompt. Let’s say we have $n$ number of test cases. Consider $\mathcal{T}$ as a set of test case functions. A test case $t_{i}\in\mathcal{T}$ is a function that maps any output response $z$ to either 00 or $1$. It can be formally defined as $t_{i}:\mathcal{Z}\rightarrow\{0,1\}$. The fitness function which is denoted as $\mathcal{F}:\mathcal{X},\mathcal{T}\rightarrow[0,1]$, maps any input prompt to a value between 00 and $1$. The fitness function is defined as follows:
$\mathcal{F}(x,\mathcal{T})=\frac{\sum_{i=1}^{i=n}t_{i}(\mathcal{S}(x))}{n}$ (2)
@@ 4.5 Baselines
To evaluate the performance of EPiC, we first identified the most relevant papers in code generation that utilized a common benchmark for comparison, i.e., papers that used HumanEval or MBPP, either exclusively or as part of their benchmarks. These two benchmarks were chosen due to their popularity in the field. Further narrowing down the selection, we applied three additional criteria: (1) whether their results were competitive (more than %90 pass@1 on HumanEval), (2) whether their GitHub repository is available online, and (3) their cost is not disproportionately high. The third criterion was specifically defined to exclude SOTA tools with excessively high costs, ensuring that the selected tools remained competitive in terms of cost-effectiveness. For comparison, we used the pass@1 metric. As a result, four papers are selected as our baselines, i.e., Reflexion[20], LATS[21], LDB[23], and AgentCoder[22] (details are in Section 2). For each baseline, we ran the experiments with the recommended configuration provided in its GitHub repository or provided in its paper. After initial evaluation, AgentCoder[22] was excluded from further analysis as we found out that it’s 6 times more expensive than the second most expensive alternative. The baselines’ configuration details used in our experiments are as follows:
Reflexion: We ran Reflexion with $max\_iters$ set to 2, $strategy$ set to “reflexion”, $pass\_at\_k$ set to 1, and $model$ set to “gpt4o”. Note that, we have adapted Reflexion’s code base to support the “gpt4o” model and tracking of time cost as these features were unavailable in Reflexion’s original version. During our experiments, we found inconsistencies between the HumanEval dataset on their GitHub and the original dataset. There were missing instances and modified or removed test cases compared to the original dataset. We reported this issue in their repository [^3] Link to Reflexion dataset issue. To ensure a valid assessment of their work, we used the original HumanEval dataset.
LATS: We ran the experiments with $max\_iters$ set to eight, $number\_of\_tests$ set to two, $strategy$ set to “mctc”, $language$ set to “py”, $pass@k$ set to 1, and $model$ set to “gpt4o”. LATS used the same code base as Reflexion, resulting in the same issues with the HumanEval dataset. Similar to Reflexion, there were missing instances and modified or removed test cases compared to the original dataset. We also implemented the “gpt4o” model and tracking of time cost features in its codebase. Note that, we have also identified and fixed an issue in LATS’s code related to counting the number of successful instances. In its codebase, $success$, which is used to calculate accuracy ($pass@k$), was incremented when the generated code passed the generated test cases. However, $success$ should be defined as passing the original test cases, not the generated ones. We corrected this issue in their code (and reported it to the authors) before running the experiments. While they have now resolved this issue in their codebase yet, the results in their paper had not been updated at the time of writing this paper [^4] LATS bug link. Like Reflexion, they also used LLM-generated test cases for internal evaluations.
LDB: We set $pass@k$ to $1$, $n\_proc$ to $1$, $max\_iters$ to $10$, $strategy$ to “ldb”, and the model to “gpt4o”. Consistent with previous works, we implemented the “gpt4o” model and tracking of time cost features as they were unavailable in the original version. The authors used seed programs in their code to enhance their results. The seed programs served as the initial implementation of the code. We removed these seed files to ensure fairness, as other baselines do not use any seeds (an initial given output code) to boost their results. Upon reviewing their dataset, we found no mismatches with the original dataset. However, we identified inconsistencies between the actual test cases in the function descriptions and the test cases they used as their visible test cases in the codebase, in HumanEval dataset. It appears that some test cases were manually modified. This poses an issue for a fair comparison. In instances where we identified such cases, we adjusted them to ensure no manual intervention. As a result, we replaced the original extracted test cases with the visible test cases to ensure a fair comparison before conducting our experiments and we reported this issue in their GitHub repository to be fixed [^5] LDB GitHub issue link.
For the MBPP experiments, LATS and Reflexion employed the same code generation approach to produce test cases, using these for internal evaluations. In contrast, LDB utilized the test cases within the MBPP dataset for the internal evaluation. We determined that using the original test cases for internal evaluation is unfair to other baselines and impractical, as these test cases are meant solely for final evaluation, not for use during the process (in practice, test cases are not typically available before the code is developed). Consequently, we modified their test cases to use the generated test cases produced by the method employed in LATS and Reflexion, so we can have a practical approach implemented across the baselines and a fair comparison.
4.6 Experimental Setup
TABLE II: Cost and effectiveness of LDB, LATS, Reflexion, and EPiC on HumanEval and MBPP datasets. The cells with green background color indicate the best scores and the red ones indicate the worst scores. EPiC results are the mean over 10 runs.
$Dataset$ $Tool$ $pass@1$ $cost$ $time(minutes)$ $mean$ $cost$ $per$ $instance$ $reported$ $pass@1$

### v2

@@ 4.3. Evaluation Metrics
To evaluate the quality of the code generated by LLMs, we utilized the pass@k metric. This metric, introduced in (Chen et al., 2021), is advantageous over metrics such as CodeBLEU and ROUGE because it evaluates code behavior rather than text semantics, which are the focus of the latter metrics. The formula for pass@k is presented in Equation 1. In this equation, $E$ denotes the expected value over the problems, $n$ is the total number of generated samples per task (code samples), and $c$ is the number of correct samples. For a single problem, this metric estimates the probability that at least one of the top $k$ samples is correct. When applied to multiple problems, it evaluates the expectation across all problems.
@@
4.3.2. ATSP
To measure the cost-effectiveness, we compare the token usage of a code-generation agent against a baseline approach. Specifically, we define a new metric, Additional Token usage per Solved Problem (ATSP), which quantifies the additional token consumption required to solve more problems compared to the baseline. For the baseline, we use a zero-shot prompting approach. Given a code generation agent $m$ and a baseline method $b$, the ATSP metric is formally defined as:
$\text{ATSP}(m,b)=\frac{T_{m}-T_{b}}{\left(P_{m}-P_{b}\right)\times N}$
where:
  • $T_{m},T_{b}$: Total token usage for method $m$ and baseline $b$, respectively
  • $P_{m},P_{b}$: pass@1 scores (fraction of problems solved) for method $m$ and baseline $b$, respectively
  • $N$: Total number of evaluation problems
The numerator ($T_{m}-T_{b}$) represents the token cost difference, while the denominator ($(P_{m}-P_{b})\times N$) reflects the absolute increase in solved problems. Lower ATSP values indicate better cost-effectiveness, as fewer additional tokens are consumed per additional solved problem compared to the baseline. Note: ATSP is only meaningful when $P_{m}>P_{b}$; methods failing to outperform the baseline are trivially less cost-effective.
4.4. Objective Function
We use $\mathcal{X}$ and $\mathcal{Z}$ to denote the prompt space and output responses respectively. Let $S:\mathcal{X}\rightarrow\mathcal{Z}$ be the LLM that takes an input $x\in\mathcal{X}$ and outputs a response $z\in\mathcal{Z}$. The output response is the code implementation of the given prompt. Let’s say we have $n$ number of test cases. Consider $\mathcal{T}$ as a set of test case functions. A test case $t_{i}\in\mathcal{T}$ is a function that maps any output response $z$ to either 0 or 111. It can be formally defined as $t_{i}:\mathcal{Z}\rightarrow\{0,1\}$. The fitness function which is denoted as $\mathcal{F}:\mathcal{X},\mathcal{T}\rightarrow[0,1]$, maps any input prompt to a value between 0 and 111. The fitness function is defined as follows:
(2) $\mathcal{F}(x,\mathcal{T})=\frac{\sum_{i=1}^{i=n}t_{i}(\mathcal{S}(x))}{n}$
@@ 4.5. Baselines
To evaluate the performance of EPiC, we first identified the most relevant papers in code generation that utilized a common benchmark for comparison, i.e., papers that used HumanEval or MBPP, either exclusively or as part of their benchmarks. These two benchmarks were chosen due to their popularity in the field. Further narrowing down the selection, we applied three additional criteria: (1) whether their results were competitive (more than 90% pass@1 on HumanEval), (2) whether their GitHub repository is available online, and (3) their cost is not disproportionately high. The third criterion was specifically defined to exclude SOTA tools with excessively high costs, ensuring that the selected tools remained competitive in terms of cost-effectiveness. For comparison, we used the pass@1 metric. As a result, four papers are selected as our baselines, i.e., Reflexion (Shinn et al., 2023), LATS (Zhou et al., 2023), LDB (Zhong et al., 2024), and AgentCoder (Huang et al., 2023b) (details are in Section 2). For each baseline, we ran the experiments with the recommended configuration provided in its GitHub repository or provided in its paper. After the initial evaluation, AgentCoder (Huang et al., 2023b) was excluded from further analysis as we found that it is 6 times more expensive than the second most expensive alternative. The baseline configurations used in our experiments are as follows:
Reflexion (Shinn et al., 2023): We ran Reflexion with $max\_iters$ set to 2, $strategy$ set to “reflexion”, $pass\_at\_k$ set to 1, and $model$ set to “o3-mini”. Note that we have adapted Reflexion’s code base to support the “o3-mini” model and tracking of token usage as these features were unavailable in Reflexion’s original version.
LATS (Zhou et al., 2023): We ran the experiments with $max\_iters$ set to eight, $number\_of\_tests$ set to two, $strategy$ set to “mctc”, $language$ set to “py”, $pass@k$ set to 1, and $model$ set to “o3-mini”. We also implemented the “o3-mini” model and tracking of token usage features in its codebase.
LDB (Zhong et al., 2024): We set $pass@k$ to 111, $n\_proc$ to 111, $max\_iters$ to 101010, $strategy$ to “ldb”, and the model to “o3-mini”. Consistent with previous works, we implemented the “o3-mini” model and tracking of token usage features as they were unavailable in the original version. The authors used Reflexion’s generated solutions as seed programs to enhance their results. The seed programs served as the initial implementation of the code. We removed these seed files to ensure fairness, as other baselines do not use any seeds (an initial solution) to boost the performance. LDB uses the provided ground truth tests in the dataset for internal evaluation and feedback to the LLM. To enable a fairer comparison among the agents, we adopted a consistent test generation approach to produce test cases rather than using the ground-truth tests for internal evaluation. Using the original test cases for internal evaluation would be unfair to other baselines and impractical, as these test cases are meant solely for final evaluation, and are not typically available before code development. Consequently, we modified their test cases to use the generated test cases produced by the method employed in LATS and Reflexion (and ours).
4.6. Experimental Setup
Table 1. Default configuration parameters for EPiC
$mutation\_probability$ $population\_size$ $top\_n$ $mutation\_alg$ $temperature$
0.40.40.4 555 101010 $sim\_words\_as\_mutator$ 0.00.00.0
For the first RQ, we used HumanEval+, MBPP+, and BigCodeBenchHard. We used O3-mini, Sonnet 3.7, and Deepseek-v3 as the backend LLM for this RQ. For the second RQ, we used the same three datasets and O3-mini as the backend LLM. For the ablation study in Section 6, we used BigCodeBenchHard dataset and o3-mini as the LLM. We used o3-mini for RQ2 and Section 6 as we found it to be the best-performing LLM in RQ1.
We used the default configuration setup for EPiC presented in Table 1. The Mutation_probability parameter specifies the likelihood of a word being substituted with a similar word. The Population_size refers to the number of populations processed for each dataset instance, while Top_n indicates the top number of related words considered for the mutation process. The temperature is the LLM temperature for code and test case generation. The defaults maintain a balance between exploration and stability. A mutation probability of 0.4 introduces sufficient variation while preserving the prompt’s intent. A population size of 5 ensures diversity without being overwhelming, and considering the top 10 related words keeps substitutions semantically relevant. Using $sim\_words\_as\_mutator$ is cost-effective, and a temperature of 0.0 guarantees reliable results.
5. Experiment Results
Table 2. Performance comparison of various LLMs on HumanEval+, MBPP+, and BigCodeBenchHard datasets under Zero-shot and EPiC. The table presents pass@1 accuracy, token usage (in thousand), and associated costs for both Zero-shot and EPiC. The $\times$ symbol indicates the factor by which the token usage in EPiC exceeds that in the zero-shot setting.
Dataset LLM Zero-shot Zero-shot Token Usage EPiC EPiC Token Usage EPiC Cost
(pass@1) (thousand tokens) (pass@1) (thousand tokens) (\$)
HumanEval+ o3-mini 0.880 170 0.945 (+6.5%) 410 (2.4×2.4$\times$2.4 ×) 1.55
claude 3.7 0.817 0.914 (+9.7%) 370 ($\times2$) 3.74
deepseek v3 0.866 0.932 (+6.6%) 540 (3.2×3.2$\times$3.2 ×) 1.26
MBPP+ o3-mini 0.680 300 0.743 (+6.3%) 1000 ($\times3$) 4.03
claude 3.7 0.674 0.724 (+5%) 800 (2.7×2.7$\times$2.7 ×) 9
deepseek v3 0.682 0.748 (+6.6%) 1040 (3.5×3.5$\times$3.5 ×) 2.53
BigCodeBenchHard o3-mini 0.350 200 0.419 (+6.9%) 1660 (8×8$\times$8 ×) 5.9
claude 3.7 0.328 0.356 (+2.8%) 1180 (6×6$\times$6 ×) 9.98
deepseek v3 0.338 0.356 (+1.8%) 1320 (7×7$\times$7 ×) 2.72
Table 3. Comparison of different approaches on HumanEval+, MBPP+, and BigCodeBenchHard datasets in terms of Pass@1 accuracy, token usage, and cost-effectiveness (ATSP). The best Pass@1 and ATSP values for each dataset are highlighted in bold.
Dataset Method Token usage Cost Performance Cost Effectiveness
(thousand tokens) (\$) (Pass@1) (ATSP)
HumanEval+ EPiC 407 1.551.551.55 0.95 20
Reflexion 422 1.541.541.54 0.92 38
LATS 2429 9.309.309.30 0.93 275
LDB 489 1.431.431.43 0.89 196
Zero-shot 167 0.660.660.66 0.88 -
MBPP+ EPiC 1003 4.034.034.03 0.74 31
Reflexion 970 3.483.483.48 0.72 44
LATS 5614 17.5817.5817.58 0.74 234
LDB 1223 3.623.623.62 0.71 81
Zero-shot 298 1.251.251.25 0.68 -
BigCodeBenchHard EPiC 1666 5.905.905.90 0.42 141
Reflexion 1023 3.023.023.02 0.37 277
LATS 6773 19.0019.0019.00 0.39 1110
LDB 2204 5.905.905.90 0.38 450
Zero-shot 202 0.870.870.87 0.35 -
5.1. RQ1: How does EPiC perform across different SOTA LLMs?
In this RQ, we evaluate EPiC’s performance in terms of functional correctness (measured by pass@1) and token usage across three benchmark datasets and using three SOTA LLMs (o3-mini, claude 3.7, and deepseek v3). The results, as summarized in Table 2, indicate that EPiC consistently outperforms the baseline prompting approach across all settings while maintaining low costs.
For example, on the HumanEval+ dataset, pass@1 increases by approximately 6%, 10%, and 7% for the respective LLMs, while token usage increases by a factor of 2.2 to 3.2 compared to zero-shot prompting. On the MBPP dataset, improvements in pass@1 range from 5% to 7%, with token usage rising by a factor of 2.7 to 3.5. In the case of BigCodeBench, pass@1 increases by 2% to 7%, and token usage increases by a factor of 6 to 8. Evidently, the more challenging the dataset, the greater the increase in token usage; for instance, HumanEval exhibits the lowest increase, whereas BigCodeBenchHard, the most challenging dataset, shows the highest token usage.
When comparing the LLMs, Claude 3.7 shows the lowest pass@1 and the highest cost, suggesting that it is less cost-effective than both o3-mini and deepseek v3. Meanwhile, o3-mini outperforms deepseek v3 on two datasets but incurs higher token costs. This pattern indicates that higher cost does not necessarily correlate with higher performance.
The results highlight a clear trade-off: EPiC achieves higher functional correctness at the cost of increased token usage. However, this trade-off should be interpreted in the context of cost-effectiveness. Even though more tokens are used per instance, the incremental token consumption (and associated monetary cost) per additional problem solved remains low. For instance, despite o3-mini’s increased token usage on HumanEval+ (from 170 to 410 thousand tokens), the overall cost remains modest (\$1.55). Similar analyses for Claude 3.7 and deepseek v3 indicate that the additional cost is justified by the improved accuracy. Nonetheless, to fully assess the cost-effectiveness of EPiC, it must be compared to SOTA code generation agents that employ iterative loops for code enhancement like EPiC. In the next research question, we will compare the cost-effectiveness of EPiC with these SOTA baselines.
Answer to RQ1: EPiC improves pass@1 across all LLMs and datasets while increasing token usage. Gains range from 2% to 10%, with token usage rising 2.2$\times$ to 8$\times$. o3-mini outperforms deepseek v3 but at higher costs, while Claude 3.7 is the least cost-effective. Despite increased token consumption, the accuracy gains justify the cost.
5.2. RQ2: How does EPiC compare to other iterative-based agents?
To evaluate EPiC’s performance relative to Reflexion, LATS, and LDB, we compare pass@1 accuracy, token usage, and cost-effectiveness across three benchmark datasets: HumanEval+, MBPP+, and BigCodeBench (see Table 3).
EPiC consistently outperforms the baselines in terms of functional correctness across all datasets. In HumanEval+, EPiC achieves a pass@1 of 0.95, surpassing Reflexion (0.92), LATS (0.93), and LDB (0.89). Similarly, in MBPP+, EPiC attains a pass@1 of 0.74, slightly outperforming Reflexion (0.72), LATS (0.74), and LDB (0.71). In BigCodeBench, EPiC achieves a pass@1 of 0.42, which is higher than Reflexion (0.37), LATS (0.39), and LDB (0.38). These results indicate that EPiC consistently generates more functionally correct solutions compared to the iterative-based code generation agents. Overall, LATS ranks second, Reflexion third, and LDB fourth in terms of pass@1 accuracy.
While EPiC demonstrates strong accuracy, it achieves this with a lower token usage than LATS and LDB, and it’s comparable to Reflexion. For HumanEval+, EPiC utilizes 407k tokens, which is less than LATS (2,429k) and LDB (490k), and only marginally less than Reflexion (422k). In MBPP+, EPiC uses 1,003k tokens, while LATS consumes 5,614k tokens, Reflexion 970k, and LDB 1,223k. For BigCodeBench, EPiC’s token usage is 1,666k, significantly lower than LATS (6,773k) and LDB (2,204k), but higher than Reflexion (1,023k). From a cost perspective, EPiC is also highly efficient. In HumanEval+, EPiC brings a total cost of \$1.55, similar to Reflexion (\$1.54) and lower than LATS (\$9.30). In MBPP+, EPiC costs \$4.03, compared to Reflexion (\$3.48), LATS (\$17.58), and LDB (\$3.62). In BigCodeBench, EPiC maintains cost-effectiveness at \$5.90, on par with LDB but significantly lower than LATS (\$19.00).
The most critical metric, however, is cost-effectiveness, measured as Incremental Token Usage per Solved Problem (ATSP), as defined in Section 4.3. This metric accounts for both cost and effectiveness by comparing token usage and pass@1 accuracy relative to a zero-shot prompting baseline. EPiC demonstrates the best cost-effectiveness (lowest ATSP score) across all datasets. For example, on HumanEval+, EPiC consumes 20k extra tokens to find an additional correct solution relative to the zero-shot baseline, whereas Reflexion requires 38k, LATS 275k, and LDB 196k tokens. Thus, EPiC is almost twice as cost-effective as Reflexion, 13 times more cost-effective than LATS, and 10 times more cost-effective than LDB. Reflexion ranks second, followed by LDB, with LATS being the least cost-effective.
A similar trend is observed for MBPP+ and BigCodeBenchHard. On MBPP+, EPiC requires 31k tokens per additional correct solution, while Reflexion uses 44k, LATS 234k, and LDB 81k, making EPiC 1.5 times more cost-effective than Reflexion, 7 times more cost-effective than LATS, and 2.5 times more cost-effective than LDB. On BigCodeBenchHard, EPiC’s ATSP is 141k, whereas Reflexion, LATS, and LDB have ATSP values of 277k, 1,110k, and 450k, respectively, further demonstrating EPiC’s better cost-effectiveness in the benchmark.
Answer to RQ2: EPiC outperforms Reflexion, LATS, and LDB in terms of functional correctness (pass@1) across all benchmark datasets. It achieves higher accuracy while maintaining lower token usage compared to LATS and LDB, and comparable effectiveness to Reflexion. In terms of cost-effectiveness, measured by Incremental Token Usage per Solved Problem (ATSP), EPiC is the most cost-effective across all datasets, requiring significantly fewer tokens to generate additional correct solutions. Reflexion ranks second, followed by LDB, with LATS being the least cost-effective.
Table 4. Ablation study on the different mutation strategy and population size on EPiC using BigCodeBenchHard dataset and o3-mini
Row Mutator Pass@1 Token Usage (thousand tokens) Effectiveness (ATSP) Population Size